# Beyond Subspace Isolation: Many-to-Many Transformer for Light Field Image Super-resolution

### v2

@@ IV Experiments
| * Geometric self-ensemble strategy is applied. | ||||||
@@ IV-E Angular Consistency
| Method | #Params. (M) | Memory (GB) | FLOPs (G) | Time (s) | PSNR/SSIM |
| LF-IINet [58] | 4.886 | 1.99 | 57.36 | 1.55 | 29.04/0.9188 |
| LFSSR-SAV [27] | 1.542 | 8.25 | 99.45 | 3.20 | 29.37/0.9223 |
| DistgSSR [22] | 3.582 | 4.43 | 65.26 | 1.89 | 28.99/0.9195 |
| HLFSR [62] | 13.865 | 2.43 | 45.73 | 6.83 | 29.20/0.9222 |
| LFT [40] | |||||
| #blocks = 4 | 1.163 | 6.41 | 30.20 | 6.22 | 29.33/0.9196 |
| #blocks = 8 | 2.150 | 11.11 | 55.64 | 12.27 | 29.44/0.9219 |
| #blocks = 12 | 3.136 | 16.80 | 81.07 | 16.80 | 29.59/0.9238 |
| EPIT [41] | |||||
| #blocks = 4 | 1.212 | 7.25 | 57.87 | 2.25 | 29.20/0.9170 |
| #blocks = 5 | 1.470 | 8.63 | 74.15 | 2.61 | 29.31/0.9196 |
| #blocks = 8 | 2.246 | 12.77 | 110.98 | 3.77 | 29.50/0.9212 |
| #blocks = 12 | 3.328 | 18.30 | 164.09 | 5.28 | 29.58/0.9212 |
| #blocks = 14 | 3.797 | 21.06 | 190.65 | 6.12 | 29.53/0.9216 |
| LF-DET [23] | |||||
| #blocks = 3 | 1.293 | 13.80 | 39.17 | 3.79 | 29.21/0.9199 |
| #blocks = 4 | 1.697 | 17.83 | 51.20 | 4.81 | 29.42/0.9220 |
| #blocks = 5 | 2.080 | 21.86 | 63.23 | 5.91 | 29.47/0.9228 |
| M2MT-Net (Ours) | |||||
| #blocks = 3 | 1.557 | 2.72 | 14.40 | 1.14 | 29.30/0.9222 |
| #blocks = 4 | 2.043 | 3.20 | 18.29 | 1.33 | 29.50/0.9226 |
| #blocks = 5 | 2.529 | 3.67 | 22.18 | 1.52 | 29.51/0.9239 |
| #blocks = 6 | 3.015 | 4.15 | 26.07 | 1.71 | 29.58/0.9253 |
| #blocks = 7 | 3.501 | 4.62 | 29.96 | 1.90 | 29.67/0.9265 |
| #blocks = 8 | 3.986 | 5.10 | 33.85 | 2.09 | 29.85/0.9284 |
@@ IV-G2 Correlation Tensor Channels
| Method | FLOPs (G) | Time (s) | #Params. (M) | Memory (GB) | PSNR/SSIM |
| 31.21 | 2.04 | 2.460 | 5.04 | 29.60/0.9256 | |
| 32.50 | 2.05 | 3.198 | 5.08 | 29.67/0.9265 | |
| 33.85 | 2.09 | 3.986 | 5.10 | 29.85/0.9284 | |
| 35.24 | 2.11 | 4.824 | 5.12 | 29.72/0.9265 |

### v3

@@ IV Experiments
* Geometric self-ensemble strategy is applied. |
@@ IV-E Angular Consistency
Method |
PSNR/SSIM | ||||||||
LF-IINet [58] | 4.886 | 1.99 | 57.36 | 1.55 | 29.04/0.9188 | ||||||||
LFSSR-SAV [27] | 1.542 | 8.25 | 99.45 | 3.20 | 29.37/0.9223 | ||||||||
DistgSSR [22] | 3.582 | 4.43 | 65.26 | 1.89 | 28.99/0.9195 | ||||||||
HLFSR [62] | 13.865 | 2.43 | 45.73 | 6.83 | 29.20/0.9222 | ||||||||
LFT [40] | |||||||||||||
#blocks = 4 | 1.163 | 6.41 | 30.20 | 6.22 | 29.33/0.9196 | ||||||||
#blocks = 8 | 2.150 | 11.11 | 55.64 | 12.27 | 29.44/0.9219 | ||||||||
#blocks = 12 | 3.136 | 16.80 | 81.07 | 16.80 | 29.59/0.9238 | ||||||||
EPIT [41] | |||||||||||||
#blocks = 4 | 1.212 | 7.25 | 57.87 | 2.25 | 29.20/0.9170 | ||||||||
#blocks = 5 | 1.470 | 8.63 | 74.15 | 2.61 | 29.31/0.9196 | ||||||||
#blocks = 8 | 2.246 | 12.77 | 110.98 | 3.77 | 29.50/0.9212 | ||||||||
#blocks = 12 | 3.328 | 18.30 | 164.09 | 5.28 | 29.58/0.9212 | ||||||||
#blocks = 14 | 3.797 | 21.06 | 190.65 | 6.12 | 29.53/0.9216 | ||||||||
LF-DET [23] | |||||||||||||
#blocks = 3 | 1.293 | 13.80 | 39.17 | 3.79 | 29.21/0.9199 | ||||||||
#blocks = 4 | 1.697 | 17.83 | 51.20 | 4.81 | 29.42/0.9220 | ||||||||
#blocks = 5 | 2.080 | 21.86 | 63.23 | 5.91 | 29.47/0.9228 | ||||||||
M2MT-Net (Ours) | |||||||||||||
#blocks = 3 | 1.557 | 2.72 | 14.40 | 1.14 | 29.30/0.9222 | ||||||||
#blocks = 4 | 2.043 | 3.20 | 18.29 | 1.33 | 29.50/0.9226 | ||||||||
#blocks = 5 | 2.529 | 3.67 | 22.18 | 1.52 | 29.51/0.9239 | ||||||||
#blocks = 6 | 3.015 | 4.15 | 26.07 | 1.71 | 29.58/0.9253 | ||||||||
#blocks = 7 | 3.501 | 4.62 | 29.96 | 1.90 | 29.67/0.9265 | ||||||||
#blocks = 8 | 3.986 | 5.10 | 33.85 | 2.09 | 29.85/0.9284 | ||||||||
@@ IV-G2 Correlation Tensor Channels
Method |
PSNR/SSIM | ||||||||
31.21 | 2.04 | 2.460 | 5.04 | 29.60/0.9256 | |||||||||
32.50 | 2.05 | 3.198 | 5.08 | 29.67/0.9265 | |||||||||
33.85 | 2.09 | 3.986 | 5.10 | 29.85/0.9284 | |||||||||
35.24 | 2.11 | 4.824 | 5.12 | 29.72/0.9265 |